# Onsager-Machlup Functional and Large Deviation Principle for Stochastic Hamiltonian Systems

### v1

@@ 5. Large Deviation Principle and Rate Function for Stochastic Hamiltonian Systems
$\begin{cases}y_{q}(t)=\varphi_{q}(t)+\epsilon\int_{0}^{t}\sigma_{q}(s)\,{\rm d}W_{q}(s),\\ y_{p}(t)=\varphi_{p}(t)+\epsilon\int_{0}^{t}\sigma_{p}(s)\,{\rm d}W_{p}(s).\end{cases}$
To simplify the notation in the proof, we introduce the term $W^{\sigma}(t):=\left(W^{\sigma}_{q}(t),W^{\sigma}_{p}(t)\right)$, which encapsulates the stochastic perturbation in the system:
$W^{\sigma}_{q}(t):=\int_{0}^{t}\sigma_{q}(s)\,{\rm d}W_{q}(s),\quad W^{\sigma}_{p}(t):=\int_{0}^{t}\sigma_{p}(s)\,{\rm d}W_{p}(s).$
Then we introduce a new probability measure $\tilde{\mathbb{P}}$, under which the transformed Brownian motions are given by:
$\tilde{W}_{q}(t)$ $=W_{q}(t)-\frac{1}{\epsilon}\int_{0}^{t}\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right)\,{\rm d}s,$
$\tilde{W}_{p}(t)$ $=W_{p}(t)-\frac{1}{\epsilon}\int_{0}^{t}\sigma_{p}^{-1}(s)\left(-\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})-\dot{\varphi}_{p}(s)\right)\,{\rm d}s.$
Under this new measure $\tilde{\mathbb{P}}$, the system is purely driven by the new Brownian motions $\tilde{W}_{q}(t)$ and $\tilde{W}_{p}(t)$, free of the deterministic drift terms.
The Radon-Nikodym derivative $\mathcal{R}:=\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}$, representing the change of measure from $\mathbb{P}$ to $\tilde{\mathbb{P}}$, is given by the exponential martingale associated with the removed drift terms. For the position variable $q$, the Radon-Nikodym derivative is:
$\frac{d\tilde{\mathbb{P}}_{q}}{d\mathbb{P}_{q}}=\exp$$\left\{\frac{1}{\epsilon}\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right),\,{\rm d}W_{q}(s)\right\rangle\right.$
$\left.-\frac{1}{2\epsilon^{2}}\int_{0}^{1}\left|\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right)\right|^{2}\,{\rm d}s\right\},$
and similarly for the momentum variable $p$:
$\frac{d\tilde{\mathbb{P}}_{p}}{d\mathbb{P}_{p}}=\exp$$\left\{\frac{1}{\epsilon}\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\left(-\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})-\dot{\varphi}_{p}(s)\right),\,{\rm d}W_{p}(s)\right\rangle\right.$
$\left.-\frac{1}{2\epsilon^{2}}\int_{0}^{1}\left|\sigma_{p}^{-1}(s)\left(\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})+\dot{\varphi}_{p}(s)\right)\right|^{2}\,{\rm d}s\right\}.$
@@
$\mathcal{R}=\exp$$\left\{\frac{1}{\epsilon}\left(\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right),\,{\rm d}W_{q}(s)\right\rangle\right.\right.$
$\qquad\left.-\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\left(\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})+\dot{\varphi}_{p}(s)\right),\,{\rm d}W_{p}(s)\right\rangle\right)$
@@
$-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}\left|\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right)\right|^{2}\,{\rm d}s\right.$
@@
$\qquad\left.\left.+\int_{0}^{1}\left|\sigma_{p}^{-1}(s)\left(\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})+\dot{\varphi}_{p}(s)\right)\right|^{2}\,{\rm d}s\right)\right\}.$
@@
(5.4) $\quad\mathbb{P}(X(t)\in K(\varphi,\epsilon))=\mathbb{P}\left(\left\|(q,p)-(\varphi_{q},\varphi_{p})\right\|\leq\epsilon\right)$
=~$=\tilde{\mathbb{P}}\left(\left\|(Y_{q},Y_{p})-(\varphi_{q},\varphi_{p})\right\|\leq\delta\right)=\mathbb{E}\left(\mathcal{R}\mathbb{I}_{\left\|W^{\sigma}\right\|\leq 1}\right)$
$=\mathbb{E}\left(\mathcal{R}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1)$
$=\exp\left\{-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}\left\|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right.\right.$
$\qquad\qquad\qquad\qquad\left.\left.+\int_{0}^{1}\left\|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right)\right\}$
×$\quad\times\mathbb{E}\left(\exp\left\{\frac{1}{\epsilon^{2}}\sum_{i=1}^{6}B_{i}\right\}\mathbb{I}_{\left\|W^{\sigma}\right\|\leq 1}\right)\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1),$
@@
$B_{1}$ $=\epsilon\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\frac{\partial H}{\partial y_{p}}(y_{q},y_{p}),\,{\rm d}W_{q}(s)\right\rangle-\epsilon\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\frac{\partial H}{\partial y_{q}}(y_{q},y_{p}),\,{\rm d}W_{p}(s)\right\rangle,$
$B_{2}$ $=-\epsilon\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\dot{\varphi}_{q}(s),\,{\rm d}W_{q}(s)\right\rangle-\epsilon\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\dot{\varphi}_{p}(s),\,{\rm d}W_{p}(s)\right\rangle,$
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left\|\sigma_{q}^{-1}(s)\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right\|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left\|\sigma_{q}^{-1}(s)\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right\|^{2}\,{\rm d}s,$
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}\left\|\sigma_{p}^{-1}(s)\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right\|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left\|\sigma_{p}^{-1}(s)\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})\right\|^{2}\,{\rm d}s,$
$B_{5}$ $=\int_{0}^{1}\left\langle\sigma_{q}^{-2}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right),\dot{\varphi}_{q}(s)\right\rangle\,{\rm d}s,$
$B_{6}$ $=-\int_{0}^{1}\left\langle\sigma_{p}^{-2}(s)\left(\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right),\dot{\varphi}_{p}(s)\right\rangle\,{\rm d}s.$
@@
$\exp$$\left\{-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}\left\|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right.\right.$
$\qquad\quad\left.\left.+\int_{0}^{1}\left\|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right)\right\}$
@@
$\mathbb{E}\left(\exp\left\{\frac{1}{\epsilon^{2}}\sum_{i=1}^{6}B_{i}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right\}\right),$
@@
Specifically, the deterministic component represents the behavior along the deterministic trajectory, while the correction component accounts for deviations induced by random perturbations. The small ball probabilities for Brownian motion represent fixed values that are intrinsically linked to the diffusion coefficients $\sigma_{q}$, $\sigma_{p}$, and the duration of time, yet they are independent of the variable $\varphi$. Furthermore, for convenience, we introduce the notation $U(\varphi)$ defined as:
$U(\varphi):=\left\|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}+\left\|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}.$
Large deviation theory focuses on large-scale deviations, and in this context, we are particularly concerned with the $\frac{1}{\epsilon^{2}}$ scale. Given the boundedness of $\sigma_{q,i}^{-1}(s)$ and $\sigma_{p,i}^{-1}(s)$, along with the fact that $H\in C^{3}_{b}(\mathbb{R}^{n}\times\mathbb{R}^{n},\mathbb{R})$, we can infer that
(5.5) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{1}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1,$

### v2

@@ 4 Most Probable Path
$\sigma(\theta,I):=\begin{pmatrix}\sigma_{\theta}(\theta,I)\!\!&\!\!0\\ 0\!\!&\!\!\sigma_{I}(\theta,I)\end{pmatrix}=\begin{pmatrix}\sigma_{\theta,1}&0&\cdots&0&0&0&\cdots&0\\ 0&\sigma_{\theta,2}&\cdots&0&0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\sigma_{\theta,n}&0&0&\cdots&0\\ 0&0&\cdots&0&\sigma_{I,1}&0&\cdots&0\\ 0&0&\cdots&0&0&\sigma_{I,2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0&0&0&\cdots&\sigma_{I,n}\\ \end{pmatrix}$,
and
$\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}=\begin{pmatrix}\frac{\partial\sigma_{\theta,1}}{\partial\theta_{1}}&\frac{\partial\sigma_{\theta,2}}{\partial\theta_{2}}&\cdots&\frac{\partial\sigma_{\theta,n}}{\partial\theta_{n}}&\frac{\partial\sigma_{I,1}}{\partial I_{1}}&\frac{\partial\sigma_{I,2}}{\partial I_{2}}&\cdots&\frac{\partial\sigma_{I,n}}{\partial I_{n}}\end{pmatrix}^{T}.$
@@
$\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}=\begin{pmatrix}\sigma_{\theta,1}\frac{\partial\sigma_{\theta,1}}{\partial\theta_{1}}&\sigma_{\theta,2}\frac{\partial\sigma_{\theta,2}}{\partial\theta_{2}}&\cdots&\sigma_{\theta,n}\frac{\partial\sigma_{\theta,n}}{\partial\theta_{n}}&\sigma_{I,1}\frac{\partial\sigma_{I,1}}{\partial I_{1}}&\sigma_{I,2}\frac{\partial\sigma_{I,2}}{\partial I_{2}}&\cdots&\sigma_{I,n}\frac{\partial\sigma_{I,n}}{\partial I_{n}}\end{pmatrix}^{T}.$
Based on the above relations, it is straightforward to verify that the resulting system can be expressed in a Hamiltonian structure. The Hamiltonian can be defined as
$H_{\sigma}(\varphi_{t}):=\frac{1}{2}\sum_{i=1}^{n}\left(\sigma_{\theta,i}(\theta,I)^{2}-\sigma_{I,i}(\theta,I)^{2}\right).$
Remark 3.
[…]
$\sum_{k}\sigma_{ik}\,\frac{\partial}{\partial\varphi_{j}}\Bigg{(}\sum_{m}\frac{\partial\sigma_{km}}{\partial\varphi_{m}}\Bigg{)}=\sum_{k}\sigma_{jk}\,\frac{\partial}{\partial\varphi_{i}}\Bigg{(}\sum_{m}\frac{\partial\sigma_{km}}{\partial\varphi_{m}}\Bigg{)}$
If this condition holds, it can be shown that the term $\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}$ can be formulated in the Hamiltonian structure. However, this stringent condition is intricate and challenging to verify in practice. Thus, it is not adopted in this work. In contrast, Condition $(C3)$ provides a more practical and tractable framework, making it a particularly favorable choice for analysis.
Corollary 4.3.
We consider here a special case, namely a stochastic Hamiltonian system with additive noise, whose dynamics are governed by the following stochastic differential equation:
$\mathrm{d}X(t)=J\nabla H(X(t))\,\mathrm{d}t+\sigma(t)\,\mathrm{d}W(t),$
where the diffusion coefficient $\sigma(t)$ depends only on the time variable $t$ and is independent of the state variable. According to Theorem 3.1, the Onsager-Machlup functional corresponding to this system can be derived as
$OM(\varphi,\dot{\varphi})$ $=\int_{0}^{1}\left|\sigma^{-1}(s)\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s.$
Moreover, by applying the Euler–Lagrange equation, we obtain that the most probable path $X^{*}(t)$ satisfies
$\mathrm{d}{X^{*}}(t)=J\nabla H(X^{*}(t))\,\mathrm{d}t.$
This result shows that, in the case of additive noise, the most probable path coincides exactly with the trajectory of the corresponding deterministic Hamiltonian system, without any additional drift correction induced by the noise. This behavior is in sharp contrast to the case of multiplicative noise, where stochastic perturbations generally introduce nontrivial modifications to the drift. Therefore, for a classical Hamiltonian system, the presence of additive noise does not alter the evolution of its most probable trajectory.
Corollary 4.4.
We consider a completely integrable stochastic Hamiltonian system, which is expressed as follows:
$\begin{cases}\,{\rm d}\theta_{i}(t)=\frac{\partial H_{0}}{\partial I_{i}}(\theta,I)\,{\rm d}t+\sum_{k=1}^{n}\frac{\partial H_{k}}{\partial I_{i}}(\theta,I)\,{\rm d}W_{k}(t),\\ \,{\rm d}I_{i}(t)=-\frac{\partial H_{0}}{\partial\theta_{i}}(\theta,I)\,{\rm d}t-\sum_{k=1}^{n}\frac{\partial H_{k}}{\partial\theta_{i}}(\theta_{i},I)\,{\rm d}W_{k}(t),\end{cases}$ (23)
where $\{H_{k}\}_{k=0}^{n}$ denotes a family of completely integrable Hamiltonian functions. The quantities $\theta_{i}(t)$ and $I_{i}(t)$ are, respectively, the $i$-th components of the angle vector $\theta(t)$ and the action vector $I(t)$, for $1\leq i\leq n$; $W_{k}(t)$ is independent one-dimensional standard Brownian motion for $1\leq k\leq n$. Consequently, system (23) can be written equivalently as
$\mathrm{d}X(t)=J\nabla H_{0}(X(t))\,\mathrm{d}t+J\nabla\bar{H}(X(t))\,\mathrm{d}W(t),$ (24)
where $J$ denotes the standard symplectic matrix, and $\nabla H_{0}(X)$ is the gradient of the Hamiltonian $H_{0}$ with respect to $X=(\theta,I)$. The notation $\nabla\bar{H}$ refers to the $2n\times n$ matrix whose columns are the gradients of the Hamiltonians $H_{1},H_{2},\dots,H_{n}$, namely,
$\nabla\bar{H}:=\left(\nabla H_{1}(X),\,\nabla H_{2}(X),\,\dots,\,\nabla H_{n}(X)\right).$
The process $W(t)$ is a $n$-dimensional standard Brownian motion, and all these Brownian motions are mutually independent. Due to the complete integrability of the system, for all $0\leq k\leq n$, we have $\frac{\partial H_{k}}{\partial\theta}(\theta,I)=0$. It is readily observed that condition $(C2)$ is no longer valid. However, this does not affect our conclusions, as equation (23) can be reformulated as follows:
$\begin{cases}\,{\rm d}\theta_{i}(t)=\frac{\partial H_{0}}{\partial I_{i}}(I)\,{\rm d}t+\sum_{k=1}^{n}\frac{\partial H_{k}}{\partial I_{i}}(I)\,{\rm d}W_{k}(t),\\ \,{\rm d}I_{i}(t)=0.\end{cases}$ (25)
Since the equation $\mathrm{d}I_{i}(t)=0$ is deterministic, we only need to analyze $\theta_{i}(t)$ separately. We may assume that $\frac{\partial H_{k}}{\partial I_{i}}(I)\neq 0$, which is a standard assumption. This, in particular, implies that the matrix $\sigma:=\left(\frac{\partial H_{k}}{\partial I_{i}}(I)\right)_{1\leq i,k\leq n}$ is an invertible $n\times n$ matrix. Therefore, based on Theorem 3.1, we derive the Onsager-Machlup functional for system (25)
$OM(\varphi_{\theta},\dot{\varphi_{\theta}})=\int_{0}^{1}\left\|\sigma^{-1}\left(\frac{\partial H_{0}}{\partial I}(I)-\dot{\varphi_{\theta}}\right)\right\|^{2}\,{\rm d}s.$
Furthermore, by applying the Euler-Lagrange equation, we demonstrate that the most probable path $X^{*}(t)$ satisfies the deterministic Hamiltonian equation
$\mathrm{d}{X^{*}}(t)=J\nabla H_{0}(X^{*}(t))\,\mathrm{d}t.$
This significant finding provides multi-dimensional insights into integrable systems under random perturbations. First, it offers rigorous mathematical verification of Li’s averaging principle [30] from the perspective of probabilistic variational principles, revealing the dominant role of deterministic dynamics in stochastically perturbed systems. Second, it establishes a theoretical connection between the Onsager-Machlup functional and the averaged equations, bridging probabilistic methods with deterministic analysis. Finally, the proposed methodology extends the traditional averaging theory to the realm of variational principles, providing a novel analytical framework for studying more general classes of stochastically perturbed Hamiltonian systems.
Therefore, although stochastic perturbations introduce complexity and uncertainty into Hamiltonian systems, minimizing the Onsager-Machlup functional allows us to identify the dominant factors governing system behavior, namely, the evolution along the most probable path. This finding is of significant importance in both theoretical research and practical applications. To further substantiate this conclusion, we present a specific example: the one-dimensional harmonic oscillator.
Example 4.5.
Consider a one-dimensional harmonic oscillator with the Hamiltonian $H(q,p)=\frac{p^{2}}{2m}+\frac{1}{2}kq^{2}$, where $\frac{p^{2}}{2m}$ represents the kinetic energy term, and $V(q)=\frac{1}{2}kq^{2}$ is the potential energy term. Here, $m=1$ denotes the mass, $k=1$ denotes the spring constant, and $q$ and $p$ respectively represent position and momentum. The classical Hamiltonian equations are given by
$\begin{cases}\begin{aligned} dq(t)&=p\,{\rm d}t,\quad&q(0)&=100,\\ dp(t)&=-q\,{\rm d}t,\quad&p(0)&=100.\end{aligned}\end{cases}$ (26)
$\begin{cases}\begin{aligned} dq(t)&=p\,{\rm d}t+(3+\sin(t)+\cos(\sqrt{2}(q+p)))\,{\rm d}W_{q}(t),&q(0)&=100,\\ dp(t)&=-q\,{\rm d}t+(3+\cos(t)+\sin(\sqrt{3}(q+p)))\,{\rm d}W_{p}(t),&p(0)&=100,\end{aligned}\end{cases}$ (27)
The system above clearly satisfies the conditions $(C1)$ and $(C2)$. Therefore, based on Theorem 3.1, we can derive the corresponding Onsager-Machlup functional, and, using the proofs provided earlier in this section, obtain the solution $X^{*}(t)$ for the most probable path of this stochastic Hamiltonian system, which satisfies the following equation:
$\begin{cases}\begin{aligned} dq(t)&=\left(p-2\sqrt{2}\sin(\sqrt{2}(q+p))(3+\sin(t)+\cos(\sqrt{2}(q+p)))\right)\,{\rm d}t,&q(0)&=100,\\ dp(t)&=-\left(q-2\sqrt{3}\cos(\sqrt{3}(q+p))(3+\cos(t)+\sin(\sqrt{3}(q+p)))\right)\,{\rm d}t,&p(0)&=100.\end{aligned}\end{cases}$ (28)
In Figure 1, we clearly observe the distinct contrast between the phase space trajectories of the classical harmonic oscillator, the stochastically perturbed oscillator, and the most probable phase space trajectory. In the absence of stochastic perturbations, the classical harmonic oscillator follows a stable, closed circular orbit, where momentum $p$ and position $q$ oscillate periodically, reflecting energy conservation. This circular trajectory signifies the steady exchange between kinetic and potential energy, thereby maintaining a constant Hamiltonian.
Figure 1: Comparison of phase-space trajectories among the deterministic (black bold solid line), stochastically perturbed (gray thin dashed lines), and most probable (red bold solid line) harmonic oscillators. The system is modeled with the following parameters: mass $m=1$, spring constant $k=1$, time step $dt=0.0001$, and total simulation time $T=500$. The trajectory of the classical system in phase space follows a stable circular orbit, indicating energy conservation and periodic motion. In contrast, the stochastic system incorporates multiplicative noise perturbations through the Euler-Maruyama method, resulting in trajectories that gradually diffuse away from the original orbit, exhibiting both dispersion and irregularity over time. The most probable trajectory, derived from the Onsager-Machlup functional, is located in the densest region of the stochastic trajectories, aligning with the theoretical predictions presented in this study.
However, when multiplicative noise perturbations are introduced, the trajectory deviates from this ideal path, exhibiting diffusion and irregularity. As illustrated in Figure 1, the stochastic oscillator’s trajectory gradually drifts away from the stable orbit, with random fluctuations disrupting the original pattern. This diffusion phenomenon arises because the noise induces energy fluctuations, causing deviations in $p$ and $q$, preventing the system from strictly adhering to the classical oscillator’s trajectory. Although some periodic behavior persists, the perturbations induce instability, leading to a trajectory that expands outward in phase space over time.
Despite these fluctuations, we also observe that the stochastic oscillator’s phase space trajectories remain densely concentrated near the most probable phase space trajectory. Notably, the most probable phase space trajectory remains closely aligned with the deterministic oscillator’s path, albeit with slight deviations caused by the multiplicative noise. Consequently, the most probable trajectory is not strictly conservative. This observation is consistent with our theoretical Theorems 3.1 and Section 4, further validating the conclusion that, even under stochastic perturbations, the system tends to evolve near the deterministic solution, albeit with slight deformations.
5 Large Deviation Principle and Rate Function for Stochastic Hamiltonian Systems
Since our objective is to study the behavior of the trajectories in the stochastic Hamiltonian system as the noise term approaches zero, we simplify equation (2) into the following form for convenience:
$\mathrm{d}X(t)=J\nabla H(X(t))\,\mathrm{d}t+\epsilon\sigma(t,X(t))\,\mathrm{d}W(t),$ (29)
where $J$ is the standard symplectic matrix, and $\nabla H(X)$ denotes the gradient of the Hamiltonian with respect to $X=(\theta,I)$. The matrix $\sigma(t,X)$ is the $2n\times 2n$ diffusion matrix, and $W(t)$ is the $2n$-dimensional standard Brownian motion. The parameter $\epsilon$ represents the noise intensity. Our primary focus is on analyzing the statistical behavior of this system as $\epsilon\to 0$.
Theorem 5.1.
Assuming that Conditions $(C1)$ and $(C2)$ hold, let $X^{\epsilon}(t):=(\theta(t),I(t))$ be the solution to the following stochastic Hamiltonian system (29). As $\epsilon\to 0$, the most probable path $\varphi^{*}(t)=(\varphi^{*}_{\theta}(t),\varphi^{*}_{I}(t))$ is given by the deterministic Hamiltonian equations
@@ 5 Large Deviation Principle and Rate Function for Stochastic Hamiltonian Systems
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\epsilon^{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$
For any path $X^{\epsilon}(t)$, the probability that the system deviates from the most probable path satisfies the large deviation principle
$\epsilon^{2}\ln\mathbb{P}(X^{\epsilon}(t)\in A)\approx-\inf_{\varphi\in A}S(\varphi),$ (30)
where $\varphi\in A$ denotes an arbitrarily continuous function, and $A\subset\mathbb{R}^{d}$ denotes an arbitrary measurable set, and suppose that $\varphi^{*}(t)\notin A$. Furthermore, the rate function $S(\varphi)$ is given by
[…]
with $\sigma^{-1}(t,\varphi_{t})$ being the inverses of the diffusion matrices $\sigma(t,\varphi_{t})$.
Proof.
[…]
$\mathrm{d}y(t)=\varphi(t)+\epsilon\int_{0}^{t}\sigma(s,y_{s})\,\mathrm{d}W(s).$
To simplify the notation in the proof, we introduce the term $W^{\sigma}(t)$, which encapsulates the stochastic perturbation in the system
$W^{\sigma}(t):=\int_{0}^{t}\sigma(s,y_{s})\,\mathrm{d}W(s).$
Then we introduce a new probability measure $\tilde{\mathbb{P}}$, under which the transformed Brownian motions are given by
$\tilde{W}(t)$ $=W(t)-\frac{1}{\epsilon}\int_{0}^{t}\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi_{s}}\right)\,{\rm d}s.$
Under this new measure $\tilde{\mathbb{P}}$, the system is purely driven by the new Brownian motion $\tilde{W}(t)$, free of the deterministic drift terms.
The Radon-Nikodym derivative $\mathcal{R}:=\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}$, representing the change of measure from $\mathbb{P}$ to $\tilde{\mathbb{P}}$, is given by the exponential martingale associated with the removed drift terms. This can be written in the following form:
$\mathcal{R}=\exp\left\{\frac{1}{\epsilon}\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi_{s}}\right),\,{\rm d}W(s)\right\rangle-\frac{1}{2\epsilon^{2}}\int_{0}^{1}\left|\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi_{s}}\right)\right|^{2}\,{\rm d}s\right\}.$
@@
$\quad\mathbb{P}(X(t)\in K(\varphi,\epsilon))=\mathbb{P}\left(\left\|X(t)-\varphi(t)\right\|\leq\epsilon\right)=\tilde{\mathbb{P}}\left(\left\|y(t)-\varphi(t)\right\|\leq\epsilon\right)$ (32)
$=\mathbb{E}\left(\mathcal{R}\mathbb{I}_{\left\|W^{\sigma}\right\|\leq 1}\right)=\mathbb{E}\left(\mathcal{R}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1)$
$=\exp\left\{-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}\,{\rm d}t\right)\right\}\times\mathbb{E}\left(\exp\left\{\frac{1}{\epsilon^{2}}\sum_{i=1}^{4}B_{i}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)\times\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1),$
@@
$B_{1}$ $=\epsilon\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})J\nabla H(y_{s}),\,{\rm d}W_{s}\right\rangle,$
$B_{2}$ $=-\epsilon\int_{0}^{1}\left\langle\sigma_{\theta}^{-1}(s,y_{s})\dot{\varphi}_{s},\,{\rm d}W_{s}\right\rangle,$
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s,$
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s.$
@@
$\exp\left\{-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}\,{\rm d}s\right)\right\}=\exp\left\{-\frac{1}{2\epsilon^{2}}\left(\int_{0}^{1}U(\varphi)\,{\rm d}s\right)\right\}$
@@
[…]
@@
Specifically, the deterministic component represents the behavior along the deterministic trajectory, while the correction component accounts for deviations induced by random perturbations. Large deviation theory focuses on large-scale deviations, and in this context, we are particularly concerned with the $\frac{1}{\epsilon^{2}}$ scale. Given the boundedness of $\sigma^{-1}(s,y)$, along with the fact that $H\in C^{3}_{b}(\mathbb{R}^{n}\times\mathbb{R}^{n},\mathbb{R})$, we can infer that
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{1}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1,$ (33)